# Jet launching from the Kerr black hole magnetosphere: An electrogeodesic approach

### v1

@@ Introduction
This formula yields $u^{\mu}(e_{0})_{\mu}]_{\infty}=\varepsilon$. Hence we recover that the conserved quantity $\varepsilon$ is interpretable as the particle’s specific energy measured by an inertial observer at infinity and at rest with respect to the black hole (such as the asymptotic Carter observer). For the case $\ell=-\kappa{\mathcal{P}}$ that we considering [cf. Eq. (4.15)], we may write $\ell=a\beta$ [cf. Eq. (4.18)] and a direct computation then gives
$z=\frac{1}{\sqrt{\Sigma\Delta}}\left[r_{\text{em}}^{2}+a^{2}\left(1-\frac{\beta}{\varepsilon}\right)-\sqrt{(r_{\text{em}}^{2}+a^{2}\cos^{2}{\theta})(r_{\text{em}}^{2}-2Mr_{\text{em}}+a^{2})}.\right]$ (4.35)
@@
First, fixing the spin of the black hole $a/M$, we observe that the extent of the region from which ejected particles appear blueshifted at infinity, i.e. $z<0$, grows with the ratio $\beta/\varepsilon$. For a slowly rotating black hole, i.e. $a=0.1M$ (top of Fig. 3), ejected particles are blueshifted only for very high values of the parameter $\beta/\varepsilon$. For higher value of the spin (bottom of Fig. 3 and Fig. 4), the required values of the parameter $\beta/\varepsilon$ for having a region of blueshift are much smaller. This behavior demonstrates the combined roles of both the spin and the magnetization of the black hole in accelerating charged particles for asymptotic observers.
@@ Discussion
In this work, we have studied the ejection of charged particles in the magnetized Kerr black hole and the built-up of a relativistic jet near the poles. Contrary to most approaches, our investigation is fully analytic and provides close expressions for key quantities characterizing the jet (acceleration of the charged particles, polar equilibrium, frame-dragging effect, blueshift of the particle energy). Let us briefly summarized the main points of this analytic model.
The present model relies on the simplest Kerr magnetosphere which consists in the vacuum magnetic monopole configuration. As we reviewed in the Section 3.1, this exact solution is selected from the fundamental symmetries of the Kerr geometry and corresponds to the Penrose current (2.56) of the Kerr black hole. Therefore, it is intimately tied to the existence of a Killing-Yano tensor in this geometry. In turn, this exact solution of the source-free Maxwell equations in the Kerr geometry can also be recovered as the test field limit of the Kerr-Newman magnetosphere with a non-zero magnetic charge and a zero electric one. The key motivation for focusing on this Kerr magnetosphere model is that it belongs to the few exact solutions for which the motion of charged particles is fully separable. In Section 3, we have reviewed in detail the geometrical condition (3.30) for a magnetosphere to preserve the Carter constant which ultimately ensures the separability, and have shown that among the models of magnetosphere built from the Killing symmetries, only the Killing-Maxwell system, based on the principal tensor, and the Penrose current, based on the KY tensor, satisfy this key condition. The first two sections thus provide a general self-contained review on (i) the different models of Kerr magnetospheres based on the Killing symmetries and (ii) the necessary conditions for the electrogeodesic motion to be separable. While most of the results presented in these first two sections are known, we have also included new results. In particular, in Section 2.1.3, we have shown that besides the Wald construction, one can also use the commuting Killing vectors and the Hodge duality to build a new non-vacuum exact solution of the test Maxwell equations. This provides the solution (2.42), which describes a radially growing toroidal magnetic field surrounding the Kerr geometry. Whether this new solution can be useful at the vicinity of the hole remains to be studied. This concludes the summary of the first two sections intended to review the required material for our investigation of the motion of charged particles in the magnetized Kerr black hole. Let us now summarize the application to the jet launching mechanism.
@@
  • The form of the acceleration: We have presented the analytic expression for the $4$-velocity $u^{\mu}\partial_{\mu}$ [Eq. (4.6)] and 4-acceleration $a^{\mu}\partial_{\mu}$ [Eq. (4.7)] of the charged particles in the whole black hole exterior, which allowed us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, such that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of jet of charged particles around the poles. For ${\mathcal{P}}>0$ and negatively charged particles, $\kappa<0$, the acceleration is outwards and the particles are ejected from the North pole.
@@
  • Polar equilibrium positions: We have analyzed the stable and unstable equilibrium positions of the polar motion depending on the energy and angular momentum $(\epsilon,\ell)$ of the charged particle. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$ where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the pole $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we have identified a new condition (4.22) for the (North) rotation axis to be a stable polar position: the energy of the particle must be either smaller or greater than a threshold depending on $\beta$, i.e (4.23). Instead, trajectories violating this condition admit a stable polar position at some $\theta_{\ast}\in\;]0,\pi/2[$.
@@
  • Magnetic frame-dragging effect: We have studied the fate of the frame-dragging effect in the Kerr magnetic monopole. Using the expression of the azimuthal angular velocity, i.e. (4.27), we have shown that the magnetic monopole introduces a new magnetic frame-dragging which completely dominates the gravitational frame-dragging. Indeed, we show that the former decays as $1/r^{2}$ while the later decays as $1/r^{3}$, showing that the new magnetic frame-dragging impacts charged particles on larger scales, even in this test field approximation. Furthermore, in contrast to the gravitational frame-dragging, which forces particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their charge, i.e. on $\kappa$.
@@
  • Geometry of the acceleration zone and asymptotic blueshift of charged particles: We have further characterized the condition under which charged particles ejected from the magnetosphere reach as asymptotic observer with a higher energy. By evaluating the redshift $z$ experienced by the particle [Eq. (4.35)] we have provided an analytical condition, i.e. $z>0$, for the particle to be blueshifted at large distance. Interestingly, this condition reveals the existence of a structure, through a radius which encodes the maximal extent of the region in which emitted particles escaping to infinity will be observed as blueshifted. The maximal radius of this region grows with the spin and the magnetization of the black hole.
@@
  • Numerical values for the parameters: We have checked that our model is self-consistent for concrete realistic numerical values of the parameters. For an electron [^8] In terms of SI units, one has actually $\kappa=\sqrt{\mu_{0}/(4\pi)}\,q/m$, so that, for an electron, $\kappa_{\rm e}=-5.56\;10^{7}\;{\rm m}^{1/2}{\rm kg}^{-1/2}$, which yields $\kappa_{\rm e}=-2.04\;10^{21}$ in geometrical units., $\kappa=\kappa_{\rm e}=-2.04\;10^{21}$, while for a proton, $\kappa=\kappa_{\rm p}=1.11\;10^{18}$. These huge values reflect the large dominance of the electromagnetic interaction over the gravitational one for elementary particles. On the contrary, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$, in order for the electromagnetic field to not significantly modify the spacetime metric, which is assumed to be the Kerr metric; otherwise one would have to consider the Kerr-Newman metric. For values of $a\sim M$, this is equivalent to demand $|{\mathcal{P}}|/a\ll 1$. It follows that $\beta$ can take values of order one or even much higher. For instance, for electrons, for which $|\kappa|\sim 10^{21}$, $\beta$ can be as large as $10^{10}$ for $|{\mathcal{P}}|=10^{-10}M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
@@
To our knowledge, the present model provides the simplest analytical study for the ejection of charged particles from the Kerr black hole (test) magnetosphere. It offers a first minimal model to study the jet launching within a vacuum Kerr magnetosphere based solely on the electrogeodesic motion, thus addressing one of the key challenge of field mentioned in the introduction. There are several open interesting directions suggested by our results.
@@
The present investigation is only preliminar and only a subset of the electrogeodesics has been discussed, namely the one with $\ell=-\kappa{\mathcal{P}}$. It reveals a rich phenomenology, which needs to be studied in detail. A full classification of the different possible trajectories of charged particles will be discussed elsewhere. It would also be interesting to confront the present findings with numerical GR/PIC simulations of this magnetosphere. Finally, we point that the present model can serve as a background to perform a perturbative study of the magnetosphere and the ejection of particles.
@@ Appendix A SageMath notebooks
Some computations performed in this article have been performed by means the free Python-based mathematical software system SageMath [80, 81]. The relevant Jupyter notebooks are publicly available, from the links below:
@@
    Killing and Killing-Yano magnetospheres arround a Kerr black hole:

### v2

@@ Introduction
This formula yields $u^{\mu}(e_{0})_{\mu}]_{\infty}=\varepsilon$. Hence we recover that the conserved quantity $\varepsilon$ is interpretable as the particle’s specific energy measured by an inertial observer at infinity and at rest with respect to the black hole (such as the asymptotic Carter observer). For the case $\ell=-\kappa{\mathcal{P}}$ that we are considering [cf. Eq. (4.15)], we may write $\ell=a\beta$ [cf. Eq. (4.18)] and a direct computation then gives
$z=\frac{1}{\sqrt{\Sigma\Delta}}\left[r_{\text{em}}^{2}+a^{2}\left(1-\frac{\beta}{\varepsilon}\right)-\sqrt{(r_{\text{em}}^{2}+a^{2}\cos^{2}{\theta})(r_{\text{em}}^{2}-2Mr_{\text{em}}+a^{2})}\right].$ (4.35)
@@
First, fixing the black hole spin $a/M$, we observe that the extent of the region from which ejected particles appear blueshifted at infinity, i.e. have $z<0$, grows with the ratio $\beta/\varepsilon$. For a slowly rotating black hole, e.g. $a=0.1M$ (top of Fig. 3), ejected particles are blueshifted only for very high values of the parameter $\beta/\varepsilon$. For higher value of the spin (bottom of Fig. 3 and Fig. 4), the required values of the parameter $\beta/\varepsilon$ for having a region of blueshift are much smaller. This behavior demonstrates the combined roles of both the spin and the magnetization of the black hole in accelerating charged particles for asymptotic observers.
@@ Discussion
In this work, we have studied the ejection of charged particles around a magnetized Kerr black hole and the build-up of a relativistic jet near the poles. Contrary to most approaches, our investigation is fully analytic and provides close expressions for key quantities characterizing the jet (acceleration of the charged particles, polar equilibrium, frame-dragging effect, blueshift of the particle energy). Let us briefly summarize the main points of this analytic model.
The model relies on the simplest Kerr magnetosphere, which consists of the vacuum magnetic monopole configuration. As we reviewed in Sec. 3.1, this exact solution is selected from the fundamental symmetries of the Kerr geometry and corresponds to the Penrose current (2.54) of the Kerr black hole. Therefore, it is intimately tied to the existence of a Killing-Yano tensor in this geometry. In turn, this exact solution of the source-free Maxwell equations in the Kerr geometry can also be recovered as the test field limit of the Kerr-Newman magnetosphere with a non-zero magnetic charge and a zero electric one. The key motivation for focusing on this Kerr magnetosphere model is that it belongs to the few exact solutions for which the motion of charged particles is fully separable. In Sec. 3, we have reviewed in detail the geometrical condition (3.30) for a magnetosphere to preserve the Carter constant which ultimately ensures the separability, and have shown that among the models of magnetosphere built from the Killing symmetries, only the Killing-Maxwell system, based on the principal tensor, and the Penrose current, based on the KY tensor, satisfy this key condition. The first two sections thus provide a general self-contained review on (i) the different models of Kerr magnetospheres based on the Killing symmetries and (ii) the necessary conditions for the electrogeodesic motion to be separable. While most of the results presented in these first two sections are known, although not in this self-contained form, we have also included new results. In particular, in Sec. 2.1.3, we have shown that besides the Wald construction, one can also use the two commuting Killing vectors and the Hodge duality to build a new non-vacuum exact solution of the test Maxwell equations. This provides the new solution (2.41), which describes a radially growing toroidal magnetic field surrounding the Kerr geometry. Whether this new solution can be useful in the vicinity of the black hole remains to be studied. This concludes the summary of the two sections intended to review the required material for our investigation of the motion of charged particles in the magnetized Kerr black hole (Secs. 2 and 3). Let us now summarize the application to the jet launching mechanism.
@@
  • The form of the acceleration: We have presented the analytic expression for the $4$-velocity $u^{\mu}$ [Eq. (4.6)] and 4-acceleration $a^{\mu}$ [Eq. (4.7)] of the charged particles in the entire black hole exterior, which allowed us to contemplate several key properties of the system. Contrary to the azimuthal and polar acceleration $(a^{\theta},a^{\varphi}$), the radial acceleration $a^{r}$ is shown to be proportional to both the magnetic charge $\mathcal{P}$} (carried by the magnetosphere) and the black hole spin $a$, so that it vanishes in the non-rotating and/or non-magnetized case. This demonstrates the key role played by the presence of a spinning magnetically charged magnetosphere in the ejection mechanism. Second, the polar acceleration turns out to be maximal at the equator and vanishing at the pole. In contrast, the radial acceleration vanishes at the equator and becomes maximal at the poles, suggesting the build up of a jet of charged particles preferentially around the poles. For ${\mathcal{P}}>0$ and negatively charged particles, $\kappa<0$, the acceleration is outwards and the particles are ejected from the North pole.
@@
  • Polar equilibrium positions: We have analyzed the stable and unstable equilibrium positions of the polar motion depending on the energy and angular momentum $(\varepsilon,\ell)$ of the charged particle. The difference with the geodesic motion is best parametrized by a new parameter $\beta=-\kappa{\mathcal{P}}/a$, where $\kappa$ is the charge-to-mass ratio of the particle. First, particles can reach the rotation axis $\theta=0$ only for $\ell=a\beta$. Focusing on this family of electrogeodesics, we have identified a new condition, (4.22), for the (North) rotation axis to be a stable polar position: the energy of the particle must be either smaller or greater than a threshold depending on $\beta$, cf. Eq (4.23). This case corresponds to an idealized aligned jet. Instead, trajectories violating this condition can admit a stable polar position at some $\theta_{\ast}\in\;]0,\pi/2[$ and thus correspond to a misaligned jet of particles.
@@
  • Magnetic frame-dragging effect: We have studied the fate of the frame-dragging effect in the Kerr magnetic monopole. Using the expression of the azimuthal angular velocity for $\ell=-\kappa{\mathcal{P}}$, i.e. Eq. (4.27), we have shown that the magnetic monopole introduces a new frame-dragging, which completely dominates the gravitational frame-dragging. Indeed, the former decays as $1/r^{2}$ while the latter decays as $1/r^{3}$, showing that the magnetic frame-dragging impacts charged particles on larger scales, even in this test field approximation. Furthermore, in contrast to the gravitational frame-dragging, which forces particles to rotate in a prograde trajectory, the magnetic counterpart forces particles to have prograde or retrograde motion depending on their electric charge. We emphasize that for misaligned jets of particles, this angular velocity shall encode the precession of the charged particles, and thus of the jet, around the rotation axis of the black hole. It would be interesting to confront this new formula to the recently measured precession of the M87* jet [61] and explore to which extent this could allow one to constrain the mass and spin of the black hole.
@@
  • Geometry of the acceleration zone and asymptotic blueshift of charged particles: We have further characterized the condition under which charged particles ejected from the magnetosphere reach as asymptotic observer with a higher energy. By evaluating the redshift $z$ experienced by the particle [Eq. (4.35)] we have provided an analytical condition, i.e. $z>0$, for the particle to be blueshifted at large distances. Interestingly, this condition reveals the existence of a structure, through a radius which encodes the maximal extent of the region in which emitted particles escaping to infinity will be observed as blueshifted. The maximal radius of this region grows with the spin and the magnetization of the black hole.
@@
  • Numerical values for the parameters: Our model is self-consistent for concrete realistic numerical values of the parameters. Indeed, as we have worked in the test field approximation, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$. We have seen that for the values of the charge-to-mass ratio $\kappa$ of the electron or proton, and for a spin of order $a\sim M$, the parameter $\beta$ can take values of order one or even much higher while still ensuring that $|{\mathcal{P}}|\ll M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
@@
To our knowledge, the present model provides the simplest analytical study for the ejection of charged particles from the Kerr black hole magnetosphere. It offers a first minimal model to study the jet launching within a source-free magnetosphere based solely on the electrogeodesic motion, thus addressing one of the key challenges mentioned in the introduction. There are several open interesting directions suggested by our results.
@@
The present investigation is only preliminar; in particular, only a subset of electrogeodesics has been discussed, namely those with $\ell=-\kappa{\mathcal{P}}$. It reveals a rich phenomenology, which needs to be studied in detail. A full classification of the different possible trajectories of charged particles will be discussed elsewhere. It would also be interesting to confront the present findings with numerical GR/PIC simulations of this magnetosphere. Indeed, the present model can serve as a background to perform a perturbative study of the magnetosphere and the ejection of particles. Finally, it would be interesting to explore how the present approach can be extended to more general magnetized rotating black hole solutions of general relativity that still exhibit Killing-Yano symmetries. See for instance [82]. We are leaving these different directions for future works.
@@ Appendix A SageMath notebooks
Some computations performed in this article have been performed by means of the free Python-based mathematical software system SageMath [83, 84]. The relevant Jupyter notebooks are publicly available from the following links:
@@
    Killing and Killing-Yano magnetospheres around a Kerr black hole: